OpenStudy (anonymous):

find the domain for y=-2/3x+6 answer must be expressed in set notation. how do i do this please?

OpenStudy (dape):

Is it \[ y=-\frac{2}{3x+6}?\]

OpenStudy (dape):

Or is it \[ y=-\frac{2}{3x}+6? \]

OpenStudy (anonymous):

the second one

OpenStudy (dape):

Okay, can you figure out any value of \(x\) that won't work in the forumla?

OpenStudy (anonymous):

i divided the right side by -2/3, which gave me -4

OpenStudy (dape):

You shouldn't need to do anything to this formula to find the domain.

OpenStudy (dape):

The thing is that if you try to put \(x=0\) into the formula, you get division by 0, which is undefined and so \(x=0\) can not be in the domain.

OpenStudy (dape):

The domain is basically all values of \(x\) which you can put into the formula so it makes sense.

OpenStudy (dape):

Now, all other values of \(x\) work except 0. Do you know set notation?

OpenStudy (anonymous):

no

OpenStudy (dape):

You could write "x is some real number except zero" in set notation as \(x\in\mathbb{R}\setminus\{0\}\) so if we call the domain \(D\) we have, in set-builder notation \(D=\{x\ |\ x\in\mathbb{R}\setminus\{0\}\}\). This reads "the domain is equal to the set of x's such that x is some real number except zero". The bar reads "such that". We could also write \(D=\{x\in\mathbb{R}\ |\ x\neq0\}\), this is a better/cleaner way, it reads "the domain is equal to the set of real x's such that x is not zero".

OpenStudy (anonymous):

ok thank you